Dingdong is 15 years older than Marian. Then years ago, he was twice as old as MArian. Find their age.
Answer by stanbon(75887)

Dingdong is 15 years older than Marian. Ten years ago, he was twice as old as MArian. Find their ages.
Equations:
d = m + 15
d-10 = 2(m-10)
----
Substitute for "d" and solve for "m":
m+15 - 10 = 2m-20
---
m + 5 = 2m-20
m = 25 (Marian's age now)
-----
Solve for "d":
d = m+15
d = 25+15
d = 40 (Dingdong's age now)_
================================
Cheers,
Stan H.
==============